CH/doc/97/03

Document

PSD Resolution Notes

Author: SPGC/WDMR

Date: 26 August 1997

Circulation : Charissa Staff

This document details some very brief tests on resolution of PSD ``E''
signals comparing the signals from different amplifiers. The
possibilities for improvement are considered.
Test:
-----
3 line alpha source (near 5 MeV) at GANIL (experiments E281/E259)
using ``Nobby'' preamps and 500 um Eurysis detector (#1376) at +120 V
Bias. 1MOhm Bias resistor, 270 Ohm offset resistors. The Charissa
amplifiers used were the modified ones as detailed in report
CH/doc/97/02.

Results:
--------
     Amp                     Shaping (uS)     Resolution (KeV)
Charissa ``Old''                 3.6*              111
Charissa ``New''#1 in crate      3.6*              127
Charissa ``New''#32 in crate     3.6*              139
Ortec 571                        3                 96
Ortec 571                        6                 80
*As calculated by comparison to Ortec pulse shape
Model Results for Capacitative noise
------------------------------------
For 0 pF expect 3.7 keV noise using an Ortec amplifier at 2uS.

for 500 um detector --> capacitance = 0.21pF / mm2
=> 525pF for 50mmX50mm 2D PSD'S
@1uS shaping   noise=30eV/pF  => 15.75 keV
@2uS shaping   noise=22eV/pF  => 11.55 keV
Assume we have capacitance contributions from detector as above =
525pF. What about the cabling? Guess around 3m cable. For screened
ribbon as inside chamber at GANIL have 131 pF/m between neighbouring
conductors. For RGB178B/U as used in SPGC 41 pin cabling for ANU have
96pF/m. Guestimate contribution to capacitance from cabling =3m X
131pF/m = 393 pF. This implies Total capacitance = 918 pF. Take
30eV/pF as order of magnitude estimate gives around 30 keV resolution
due to capacitative part of noise. This compares with the same order
of magnitude for measured values above. At 6uS expect to be a factor
of 2 better for this noise contribution than at 3uS. This looks
roughly true...
For the Ortec 571 ...
@3uS 30 keV (capacitative) + 70 keV (other)
@6uS 15 keV                + 70 keV
Conclusions
-----------
These noise contributions are SMALL when considering heavy ions. Yes,
a better amplifier will do better. The Ortec amp is better able to
filter the noise in the particular environment found in these tests. A
cheaper solution is to reduce length of cables (up to a factor of 2
improvement can be gained) , investigate preamp performance and use
thicker detectors. ``Reasonable'' long time constants should be used
but extending the time constant >6uS would seem to have a minor effect
in improving noise but a major problem for peak detection.  Charissa
amplifiers with the simple pulse shaping are roughly a factor of 2
worse i.e. up to 60 to 70 keV from capacitative noise effects. Proper
detailed measurements are required.

Note
----
The capacitance of strip detector + 3m of cable is about the same as

Sheet detector + no cable. So we shouldn't be surprised that both of
these combinations give similar results at ANU!